The measure of one interior angle of a regular polygon is two times the measure of one of its exterior angles. How many sides does this polygon have?
Answer by stanbon(75887)

The measure of one interior angle of a regular polygon is two times the measure of one of its exterior angles. How many sides does this polygon have?
---------
Let the interior angle be 2x; Let the exterior angle be "x":
---
X + 2x = 180
3x = 180
x = 60 degrees (each exterior angle)
----
The sum of the exterior angles is 360 degrees
----
# of exterior angles = 360/60 = 6
----
# of sides = 6
================
Cheers,
Stan H.